# Lie algebra homology with coefficients tensor products of the adjoint representation in relative polynomial degree $2$

### v1

@@ 5. The case $r=3$
The morphism $\beta$ also identifies as the composite of $\alpha\otimes\mathrm{Id}:\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}\rightarrow\Lambda^{1}\otimes\Lambda^{3}\otimes\Lambda^{1}$ with the cyclic place permutation $\Lambda^{1}\otimes\Lambda^{3}\otimes\Lambda^{1}\stackrel{{\scriptstyle\mathrm{cyclic}}}{{\rightarrow}}\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$ (this corresponds to $\Lambda^{1}\otimes F\rightarrow F\otimes\Lambda^{1}$, where $F:=\Lambda^{3}\otimes\Lambda^{1}$.)
Lemma 5.5.
Proof.
Using Lemma 5.5, Proposition 5.2 can be restated as:
Proposition 5.6.
We thus need to understand the automorphism $-\beta$ of $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$. This identifies explicitly using Corollary 4.10:
Lemma 5.7.
The automorphism $-\beta$ of $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$ evaluated on $V$ is given for $x,y,z,s,t\in V$ by
$-\beta(x\wedge y\wedge z\otimes s\otimes t)=\frac{1}{2}\big{(}x\wedge y\wedge z\otimes t\otimes s+x\wedge y\wedge s\otimes t\otimes z-x\wedge z\wedge s\otimes t\otimes y+y\wedge z\wedge s\otimes t\otimes x\big{)}.$
To analyse the automorphism $-\beta$ by exploiting the representation theory, one can use the natural decomposition of $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$ that is given by Pieri’s rule:
(24) $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}\cong\Lambda^{5}\oplus\mathbf{S}_{(3,1^{2})}\oplus\mathbf{S}_{(2^{2},1)}\oplus\mathbf{S}_{(2,1^{3})}^{\oplus 2}.$
By Schur’s lemma, we have the following analogue of Lemma 4.3:
Lemma 5.8.
There is an isomorphism of algebras
$\mathrm{End}(\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1})$$\cong$ $\mathrm{End}(\Lambda^{5})\prod\mathrm{End}(\mathbf{S}_{(3,1^{2})})\prod\mathrm{End}(\mathbf{S}_{(2^{2},1)})\prod\mathrm{End}(\mathbf{S}_{(2,1^{3})}^{\oplus 2})$
$\cong$ $\mathbb{Q}_{(1^{5})}\prod\mathbb{Q}_{(3,1^{2})}\prod\mathbb{Q}_{(2^{2},1)}\prod M_{2}(\mathbb{Q}).$
Remark 5.9.
The fact that the decomposition (24) is not multiplicity-free is the source of complexity in the proof of Theorem 5.1. In particular, the isomorphism $\mathrm{End}(\mathbf{S}_{(2,1^{3})}^{\oplus 2})\cong M_{2}(\mathbb{Q})$ is not canonical.
We analyse the isomorphism (24) in two different ways, adapted to considering the action of $\tau\circ\alpha\otimes\mathrm{Id}$ and of $\mathrm{Id}\otimes\tau$, the two isomorphisms appearing in the construction of $\beta$ in Notation 5.3.
Then, by Corollary 4.9, one has the following, in which $(1,-1)$ denotes the diagonal matrix in $M_{2}(\mathbb{Q})$ with respect to the chosen decomposition:
Lemma 5.10.
There exists an isomorphism as in (24) with respect to which $\tau\circ\alpha\otimes\mathrm{Id}\in\mathrm{End}(\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1})$ acts via the isomorphism of Lemma 5.8 by $(1,-1,-1,(1,-1))$.
Then, analogously to Lemma 5.10, one has:
Lemma 5.11.
There exists an isomorphism as in (24) with respect to which $\mathrm{Id}\otimes\tau\in\mathrm{End}(\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1})$ acts via the isomorphism of Lemma 5.8 by $(-1,1,-1,(1,-1))$.
Since there are unique composition factors of $\Lambda^{5}$, $\mathbf{S}_{(3,1^{2})}$, and $\mathbf{S}_{(2^{2},1)}$ in $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$, Lemmas 5.10 and 5.11 imply:
Lemma 5.12.
The automorphism $-\beta$ restricts to
  1. (1)
    $\mathrm{Id}$ on the summand $\Lambda^{5}\oplus\mathbf{S}_{(3,1^{2})}$;
  2. (2)
    $-\mathrm{Id}$ on the summand $\mathbf{S}_{(2^{2},1)}$.
To prove Theorem 5.1, it remains to show that $(-\beta)^{\circ 3}$ has no non-zero fixed points for the factor $\mathbf{S}_{(2,1^{3})}^{\oplus 2}$; this restriction can be considered as an element of $\mathrm{End}(\mathbf{S}_{(2,1^{3})}^{\oplus 2})\cong M_{2}(\mathbb{Q})$ via Lemma 5.8. It is sufficient to calculate the eigenvalues of this action.
Now, the decompositions corresponding to the two composition factors $\mathbf{S}_{(2,1^{3})}$ in $\Lambda^{3}\otimes\Lambda^{1}\otimes\Lambda^{1}$ that are used in Lemmas 5.10 and 5.11 do not coincide (that they must be different is a consequence of the calculations below), so one cannot directly read off the action of $-\beta$ on the corresponding isotypical component. One can, however, deduce the determinant:
Lemma 5.13.
The automorphism $-\beta$ restricted to $\mathbf{S}_{(2,1^{3})}^{\oplus 2}$, considered as an element of $M_{2}(\mathbb{Q})$ via Lemma 5.8, has determinant $1$.
Proof.
Using Lemmas 5.10 and 5.11, one sees that the determinant of $\beta$ is $1$, using that the diagonal matrix $\mathrm{diag}(1,-1)$ has determinant $-1$. Thus $-\beta$ also has determinant $1$, since $-\mathrm{Id}=\mathrm{diag}(-1,-1)$ has determinant one. ∎
This module has dimension $20$; the dimensions of $\mathrm{sgn}_{5}$, $S_{(3,1^{2})}$, $S_{(2^{2},1)}$, and $S_{(2,1^{3})}$ are respectively $1$, $6$, $5$, and $4$. The automorphism $-\beta$ induces an automorphism of the $\mathfrak{S}_{5}$-module given by (25), and hence of its underlying $\mathbb{Q}$-vector space, $\mathbb{Q}^{\oplus 20}$.
Lemma 5.14.
The element of $M_{20}(\mathbb{Q})$ associated to $-\beta$ has trace 00.
Proof.
From this, one deduces the following counterpart of Lemma 5.13
Lemma 5.15.
The automorphism $-\beta$ restricted to $\mathbf{S}_{(2,1^{3})}^{\oplus 2}$, considered as an element of $M_{2}(\mathbb{Q})$ via Lemma 5.8, has trace $-\frac{1}{2}$.
Proof.
We first calculate the trace of the restriction of $-\beta$ (considered as an element of $M_{20}(\mathbb{Q})$ as in Lemma 5.14) to the direct summand $S_{(2,1^{3})}^{\oplus 2}$ (which has underlying vector space of dimension $8$). Using the decomposition (25) and Schur’s lemma, $-\beta$ has block diagonal form. Moreover, by Lemma 5.12, it acts as $\mathrm{Id}$ on $\mathrm{sgn}_{5}\oplus S_{(3,1^{2})}$ (which has dimension $7$) and by $-\mathrm{Id}$ on $S_{(2^{2},1)}$ (which has dimension $5$). Using the block diagonal form and Lemma 5.14, it follows that the restriction of $-\beta$ to $S_{(2,1^{3})}^{\oplus 2}$ (an element of $M_{8}(\mathbb{Q})$) has trace $-2$.
By Schur’s Lemma, this element of $M_{8}(\mathbb{Q})$ is obtained from $\mathrm{End}_{\mathfrak{S}_{5}}(S_{(2,1^{3})}^{\oplus 2})\cong M_{2}(\mathbb{Q})$ by the morphism of algebras $M_{2}(\mathbb{Q})\rightarrow M_{8}(\mathbb{Q})$ induced by the functor $-\otimes\mathbb{Q}^{\oplus 4}$ on vector spaces. (The trace is independent of choices made.) The result follows. ∎
Proof of Theorem 5.1.
By Proposition 5.6, $\ker\overline{\delta}_{3}$ is the equalizer of $\mathrm{Id}$ and $(-\beta)^{\circ 3}$. The behaviour of this on the single-multiplicity composition factors is given by Lemma 5.12. To conclude, we need to show that $\mathrm{Id}-(-\beta)^{\circ 3}$ is injective restricted to $\mathbf{S}_{(2,1^{3})}^{\oplus 2}$.
Using the trace and determinant, from Lemmas 5.13 and 5.15, one deduces that the automorphism $-\beta$ restricted to $\mathbf{S}_{(2,1^{3})}^{\oplus 2}$, considered as an element of $M_{2}(\mathbb{Q})$ via Lemma 5.8, has eigenvalues (in $\mathbb{C}$) $\frac{1}{4}(-1\pm\sqrt{15}i).$ This implies that $1$ is not an eigenvalue of $(-\beta)^{\circ 3}$, whence the result. ∎
@@ 6. The proof of Theorem 1
The purpose of this section is to prove Theorem 1, based on the analysis of $\ker\overline{\delta}_{r}$ (for $r\geq 2$) by applying Proposition 3.9. The proof will exploit the lemmas concerning Schur functors given in Section 6.1.
@@ 6.1. Preparatory lemmas
Consider the exterior power functors $\Lambda^{n}\cong\mathbf{S}_{(1^{n})}$, for $n\in\mathbb{N}$. The coproduct provides the inclusion $\Lambda^{n+1}\hookrightarrow\Lambda^{n}\otimes\Lambda^{1}$; up to non-zero scalar multiple, this is the unique non-trivial natural transformation of this form. On applying $-\otimes\Lambda^{1}$, one therefore obtains $\Lambda^{n+1}\otimes\Lambda^{1}\hookrightarrow\Lambda^{n}\otimes T^{2}$. Post-composing with $\mathrm{Id}_{\Lambda^{n}}\otimes\tau$ gives a second, distinct embedding.
@@
Forming the tensor product $\otimes\Gamma^{1}$, one obtains the inclusion
@@
Post-composing with $(\mathrm{Id}\otimes\tau)$ gives the inclusion $(\mathrm{Id}\otimes\tau)\big{(}\mathbf{S}_{(n,1^{2})}\otimes\Gamma^{1})\subset\Lambda^{3}\otimes\Gamma^{n-2}\otimes T^{2}$. The following result is a counterpart of Lemma 6.1:
@@
For $n\geq 2$, the intersection of $\mathbf{S}_{(n,1^{2})}\otimes\Gamma^{1}$ and $(\mathrm{Id}\otimes\tau)\big{(}\mathbf{S}_{(n,1^{2})}\otimes\Gamma^{1}\big{)}$ in $\Lambda^{3}\otimes\Gamma^{n-2}\otimes T^{2}$ is isomorphic to $\mathbf{S}_{(n+1,1^{2})}$.
@@
in which $\mu:\Lambda^{3}\otimes\Lambda^{1}\rightarrow\Lambda^{4}$ is the product and the top square is given by the definition of the intersection $\cap$ and the above, using the embedding (26) composed with $\Lambda^{3}\otimes\Gamma^{n-2}\otimes T^{2}\subset\Lambda^{3}\otimes T^{n}$.
The labelled map is zero by Lemma 2.8, using the characterization of $\mathsf{d}$ given by Lemma 2.7. This implies that $\bigcap$ lies in the kernel of the right hand vertical composite. Since the latter factors across $\mathsf{d}\colon\Lambda^{3}\otimes\Gamma^{n}\rightarrow\Lambda^{4}\otimes\Gamma^{n-1}$, it follows that $\bigcap$ lies in $\ker\mathsf{d}\cong\mathbf{S}_{(n+1,1^{2})}$ (using Example 2.9). To conclude, it remains to show that this is an equality (this is equivalent to establishing the non-triviality of $\bigcap$).
@@
6.2. The proof of Theorem 1
@@
The cases $r\in\{1,2,3\}$ have been established in Example 3.4, Theorem 4.18, and Theorem 5.1 respectively. Hence, by induction upon $r$, we may suppose that $r>3$ and that the result has been proved for $r-1$. It remains to prove the inductive step.
@@
To establish this claim, it is useful to visualize the statement of Proposition 4.15 by considering that $i\in\{1,\ldots,r\}$ indexes the vertices of a $r-1$-simplex, analogously to the geometric visualization proposed in Section 3.3. Then the inclusion (27) is obtained by restricting to the codimension one face given by $i\neq r$, using Proposition 4.15 for $r-1$ to identify $\ker\overline{\delta}_{r-1}$. The inclusion (3.3) is obtained similarly, this time using the face $i\neq r-1$. Using the equivariance property of $\overline{\delta}_{r}$, one obtains the stated inclusion by exploiting the transposition of tensor factors $\tau$.
@@
By Lemmas 6.1 and 6.2 for the respective terms, the latter identifies with

### v4

@@ 4.6. A lower bound for $\ker\overline{\delta}_{r}$
The case $r=3$ of this result is established explicitly in the (easy part of the) proof of Theorem 6.1. Moreover, this can be used to give an alternative proof of Proposition 4.21.
@@ 5. The proof of Theorem 1
The purpose of this section is to prove Theorem 1, based on the analysis of $\ker\overline{\delta}_{r}$ (for $r\geq 2$) by applying Proposition 3.10. The (calculational) proof of the initial step of the inductive proof is postponed until Section 6 (that proof is entirely independent of the material in this section).
The proof exploits the lemmas concerning Schur functors given in Section 5.1.
@@ 5.1. Preparatory lemmas
Consider the exterior power functors $\Lambda^{n}\cong\mathbf{S}_{(1^{n})}$, for $n\in\mathbb{N}$. The coproduct provides the inclusion $\Lambda^{n+1}\hookrightarrow\Lambda^{n}\otimes\Lambda^{1}$; up to non-zero scalar multiple, this is the unique non-trivial natural transformation of this form. On applying $\otimes\Lambda^{1}$, one therefore obtains $\Lambda^{n+1}\otimes\Lambda^{1}\hookrightarrow\Lambda^{n}\otimes T^{2}$. Post-composing with $\mathrm{Id}_{\Lambda^{n}}\otimes\tau$ gives a second, distinct embedding.
@@
Applying the functor $\otimes\Gamma^{1}$, one obtains the inclusion
@@
Post-composing with $(\mathrm{Id}\otimes\tau)$ gives the inclusion $(\mathrm{Id}\otimes\tau)\big(\mathbf{S}_{(n,1^{2})}\otimes\Gamma^{1})\subset\Lambda^{3}\otimes\Gamma^{n-2}\otimes T^{2}$.
Using the canonical inclusion $\Gamma^{n-2}\subset T^{n-2}$, we may consider the functors $\mathbf{S}_{(n,1^{2})}\otimes\Gamma^{1}$ and $(\mathrm{Id}\otimes\tau)\big(\mathbf{S}_{(n,1^{2})}\otimes\Gamma^{1}\big)$ as lying in $\Lambda^{3}\otimes T^{n}$.
The following result is a counterpart of Lemma 5.1:
@@
For $n\geq 2$, the intersection of $\mathbf{S}_{(n,1^{2})}\otimes\Gamma^{1}$ and $(\mathrm{Id}\otimes\tau)\big(\mathbf{S}_{(n,1^{2})}\otimes\Gamma^{1}\big)$ in $\Lambda^{3}\otimes T^{n}$ is isomorphic to $\mathbf{S}_{(n+1,1^{2})}$.
@@
in which $\mu:\Lambda^{3}\otimes\Lambda^{1}\rightarrow\Lambda^{4}$ is the product and the top square is given by the definition of the intersection $\cap$ and the above, using the embedding (24).
The labelled map is zero by Lemma 2.9, using the characterization of $\mathsf{d}$ given by Lemma 2.8. This implies that $\bigcap$ lies in the kernel of the right hand vertical composite. Hence, using Example 2.10, $\bigcap$ lies in $\ker\mathsf{d}\cong\mathbf{S}_{(n+1,1^{2})}$. To conclude, it remains to show that this is an equality (this is equivalent to establishing the non-triviality of $\bigcap$).
@@
5.2. The proof of Theorem 1 (assuming Theorem 6.1)
@@
The cases $r\in\{1,2\}$ have been established in Example 3.5 and Theorem 4.20 respectively. The case $r=3$ is a (somewhat technical) calculation; this has been postponed until Section 6, where it is proved as Theorem 6.1.
Hence, by induction upon $r$, we may suppose that $r>3$ and that the result has been proved for r1r-1. It remains to prove the inductive step.
@@
To establish this claim, it is useful to visualize the statement of Proposition 4.17 by considering that $i\in\{1,\ldots,r\}$ indexes the vertices of a r1r-1-simplex, analogously to the geometric visualization proposed in Section 3.3. Then the inclusion (25) is obtained by restricting to the codimension one face given by $i\neq r$, using Proposition 4.17 for r1r-1 to identify $\ker\overline{\delta}_{r-1}$. The inclusion (3.3) is obtained similarly, this time using the face $i\neq r-1$. Using the equivariance property of $\overline{\delta}_{r}$, one obtains the stated inclusion by exploiting the transposition of tensor factors $\tau$.
@@
By Lemmas 5.1 and 5.2 for the respective terms, the latter identifies with